# Continuous-time persuasion by filtering

### v1

@@ 4.2 Carbon footprint accounting rules
We focus here on the second example, introduced in Section 2.2. The parameter identification (example-general theory) and the verification of the standing assumptions in this case is provided in Appendix A.2. A direct application of the results of the general case leads to the expression of the optimal control of the representative firm
$v^{\ast}_{t}=\kappa(\widehat{X}_{t}^{*}-\ell)-\beta\widehat{X}^{\ast}_{t}-\beta\hat{\ell}(q),\quad\hat{\ell}(q):=\frac{\kappa^{2}\ell-2\gamma\bar{q}}{\kappa^{2}+2\gamma\bar{\lambda}},\quad\bar{\lambda}:=\lambda_{a}+\lambda_{q},\quad\bar{q}:=\lambda_{a}a+\lambda_{q}q.$ (4.10)
The fixed point condition on the value of the stationary quantile target $q$ is given by
$q=\hat{\ell}(q)-\epsilon\sigma_{\infty}(b)=\frac{\kappa^{2}\ell-2\gamma\bar{q}}{\kappa^{2}+2\gamma\bar{\lambda}}-\epsilon\sigma_{\infty}(b),$ (4.11)
where the stationary variance of the carbon footprint process $X^{\ast}$ is given by
This last function does not depend on $q$. Hence, we find that the quantile equilibrium value $q_{\infty}^{\ast}$ is given by
$q^{\ast}_{\infty}=\frac{\kappa^{2}\ell-2\gamma\lambda_{a}a}{\kappa^{2}+2\gamma(\bar{\lambda}+\lambda_{q})}-\epsilon\sigma_{\infty}\frac{\kappa^{2}+2\gamma\bar{\lambda}}{\kappa^{2}+2\gamma(\bar{\lambda}+\lambda_{q})}.$ (4.13)
Thus, the stationary mean $m^{\ast}_{\infty}$ of the process $X$ is given by $m^{\ast}_{\infty}=\hat{\ell}(q^{\ast}_{\infty})=q^{\ast}_{\infty}+\epsilon\sigma_{\infty}(b)$, and so
$m^{\ast}_{\infty}=\frac{\kappa^{2}\ell-2\gamma\lambda_{a}a}{\kappa^{2}+2\gamma(\bar{\lambda}+\lambda_{q})}+\epsilon\frac{2\gamma\lambda_{q}}{\kappa^{2}+2\gamma(\bar{\lambda}+\lambda_{q})}\sigma_{\infty}(b)=:m_{0}+\epsilon m_{1}\sigma_{\infty}(b).$ (4.14)
@@
$\inf_{z\in(\underline{v},\bar{v}]}G(z)+\frac{1}{2}c\big{[}(m_{0}+\epsilon m_{1}z)^{2}+z^{2}\big{]},$
where we exploit the notation $z:=\sigma_{\infty}^{2}(b)$ and where the function $G$ measures the cost to reduce the standard deviation of the stationary law of $X$. A possible form that leads to closed-form expression for the optimal reduction is
$G(z)=-\frac{1}{2}\frac{\eta}{\Delta v}\ln\frac{z-\underline{v}}{\Delta v},\quad z\in(\underline{v},\bar{v}].$
For an interior solution to exist, it is necessary and sufficient that the marginal benefit exceeds the marginal cost at $z=\bar{v}$, i.e.,
$\eta\leq c(\Delta v)^{2}\big{[}\epsilon m_{1}m_{0}+(1+\epsilon^{2}m_{1}^{2})\bar{v}\big{]}.$
If this condition is not satisfied, it is optimal to set the standard deviation to its level $\bar{v}$, i.e., to provide no information. In the other case, the interior solution is given by
$\sigma^{\ast}:=\frac{1}{2}\Big{[}\sqrt{\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}-\underline{v}\Big{)}^{2}+4\frac{\eta+\epsilon m_{0}m_{1}\underline{v}}{c\Delta v(1+\epsilon^{2}m_{1}^{2})}}-\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{v}\Big{]}.$
Figure 2 (a) and (b) display $m_{\infty}^{\ast}$ and $\sigma_{\infty}^{\ast}$ as a function of the best-in-class parameter $\epsilon$. We observe that the optimal standard deviation $\sigma^{\ast}_{\infty}$ is a decreasing function of $\epsilon$, whereas the mean of stationary carbon footprint is increasing. Indeed, the mean-field Nash equilibrium of this best-in-class game consists in increasing the average carbon footprint of the representative firm to make it less costly to achieve a below average carbon footprint reduction. In a way, it is easier to look cleaner when compared to a crowd of heavy polluters. The information provision succeeds in mitigating this greenwashing perverse effect of the good intention of a best-in-class policy. Indeed, Figure 2 (a) shows that as $\epsilon$ increases the equilibrium average carbon footprint $m_{\infty}^{\ast}$ gets closer to its lowest possible bound, i.e., the information provision is more stringent and provides more efficient results. Furthermore, the information provision succeeds in reducing the standard deviation of the carbon footprint of the representative firm. Figure 2 (c) shows the decomposition of the cost function of the Sender. The black dotted line represents the cost in the absence of information provision, and thus, the damage. When providing information, the damage (red curve) is significantly reduced and this reduction is increasing with the level of best-in class intention. Besides, this reduction is obtained at an almost constant overhead cost of information provision (difference between the blue curve giving the total cost of the Sender and the red curve giving the damage part).
As a result, if we interpret the information provision as the level of stringency of the accounting rules, our model of information provision shows that the rules should be tight in sectors where firms declare a high level of best-in-class target. The level of stringency is designed to counterbalance the Nash equilibrium that leads them to collectively increase pollution to make their target easier to reach.
@@
Figure 2: As a function of $\epsilon$ (a) $m^{\ast}_{\infty}$ (b) $\sigma_{\infty}$ (c) Damage cost with optimal information provision and without any information provision. Parameters: $\kappa=0.5$, $\ell=a=9$, $\gamma=0.1$, $c=0.5$, $\lambda_{a}=0.25$, $\lambda_{q}=0.75$.
@@ A.2 Carbon footprint accounting rules
The function in the Receiver’s objective functional is $f(x,v)=(\lambda_{q}+\lambda_{a})x^{2}-2(a\lambda_{a}+q\lambda_{q})x+\lambda_{q}q^{2}+\lambda_{a}a^{2}+\frac{1}{2\gamma}{v^{2}}$, with $\lambda_{a},\lambda_{q},\gamma>0$. Hence, we have $F_{0}=\lambda_{q}q^{2}+\lambda_{a}a^{2},F_{1}=-2a\lambda_{a}-2q\lambda_{q},F_{2}=\lambda_{q}+\lambda_{a},C_{0}=0,C_{1}=\frac{1}{2\gamma}$. The function in the Sender functional is $g(x,v)=cx^{2}$, with $c>0$.

### v2

@@ 4.2 Carbon footprint accounting rules
We focus here on the second example, introduced in Section 2.2. The parameter identification (example-general theory) and the verification of the standing assumptions in this case is provided in Appendix A.2. A direct application of the results of the general case leads to the expression of the optimal control of the representative firm:
For a given value of $b$ and a given value of $q$, the optimal dynamics of the state $X^{\ast}$ and its estimate $\widehat{X}^{\ast}$ are given by
$\mathrm{d}\widehat{X}_{t}=-\beta\Big{(}\widehat{X}_{t}-\hat{\ell}(q)\Big{)}\mathrm{d}t+bP_{b}(t)\mathrm{d}I_{t},\quad\mathrm{d}X_{t}=\big{[}\kappa(\widehat{X}_{t}-X_{t})+\beta(\hat{\ell}(q)-\hat{X}_{t})\big{]}\mathrm{d}t+\mathrm{d}W_{t}$
where
^$\hat{\ell}(q):=\frac{\ell\kappa^{2}+2\gamma\bar{q}}{\kappa^{2}+2\gamma\bar{\lambda}},\quad\bar{\lambda}:=\lambda_{a}+\lambda_{q},\quad\bar{q}:=\lambda_{a}a+\lambda_{q}q,\quad\beta:=\sqrt{\kappa^{2}+2\gamma\bar{\lambda}}$
and the optimal control is given by
$v^{\ast}_{t}=\kappa(\widehat{X}_{t}^{*}-\ell)-\beta\widehat{X}^{\ast}_{t}+\beta\hat{\ell}(q).$ (4.11)
The fixed point condition on the value of the stationary target $q$ is given by $q=\hat{\ell}(q)-\epsilon\sigma_{\infty}(b)$ where the stationary variance of the carbon footprint process $X^{\ast}$ is given by
$\sigma_{\infty}^{2}(b)=\frac{1}{2\kappa}\Big{[}1-\frac{\beta-\kappa}{\beta}\Big{(}\sqrt{1+\frac{\kappa^{2}}{b^{2}}}-\frac{\kappa}{b}\Big{)}^{2}\Big{]},$ (4.12)
which is the same as in the smart meters application of section 4.1. This last function does not depend on $q$. Hence, we find that the equilibrium value $q_{\infty}^{\ast}(b)$ is given by
$q_{\infty}^{\ast}(b):=\frac{\ell\kappa^{2}+2\gamma a\lambda_{a}}{\kappa^{2}+2\gamma\lambda_{a}}-\epsilon\frac{\kappa^{2}+2\gamma\bar{\lambda}}{\kappa^{2}+2\gamma\lambda_{a}}\sigma_{\infty}(b)$ (4.13)
Thus, the stationary mean $m^{\ast}_{\infty}(b)$ of the process $X$ is given by $m^{\ast}_{\infty}(b)=\hat{\ell}(q^{\ast}_{\infty}(b))=q^{\ast}_{\infty}(b)+\epsilon\sigma_{\infty}(b)$, and so
$m^{\ast}_{\infty}(b)=\frac{\kappa^{2}\ell+2\gamma a\lambda_{a}}{\kappa^{2}+2\gamma\lambda_{a}}-\epsilon\frac{2\gamma\lambda_{q}}{\kappa^{2}+2\gamma\lambda_{a}}\sigma_{\infty}(b)=:m_{0}-\epsilon m_{1}\sigma_{\infty}(b).$ (4.14)
@@
$\inf_{z\in(\underline{\sigma},\bar{\sigma}]}G(z)+\frac{1}{2}c\big{[}(m_{0}-\epsilon m_{1}z)^{2}+z^{2}\big{]},$
where we exploit the notation $z:=\sigma_{\infty}(b)$, $\underline{\sigma}$ and $\overline{\sigma}$ have been defined in (4.4), and where the function $G$ measures the cost to reduce the standard deviation of the stationary law of $X$. A possible form that leads to closed-form expression for the optimal reduction is
$G(z)=-\frac{\eta}{\Delta\sigma}\ln\frac{z-\underline{\sigma}}{\Delta\sigma},\quad z\in(\underline{\sigma},\bar{\sigma}],\quad\Delta\sigma=\overline{\sigma}-\underline{\sigma},$
where $\eta$ is a nonnegative parameter. This form exhibits the same type of properties as the cost function $H$ used in the smart meters application of section 4.1. It is a nonincreasing, strictly convex function that satisfies $G(\overline{\sigma})=0$ and $\lim_{z\to\underline{\sigma}}G(z)=+\infty$. As in the smart meters application, it ensures that it is infinitely costly to reach perfect truth.
For an interior solution to exist, it is necessary and sufficient that the marginal benefit exceeds the marginal cost at $z=\bar{\sigma}$, i.e.,
$\eta\leq c(\Delta\sigma)^{2}\big{[}(1+\epsilon^{2}m_{1}^{2})\bar{\sigma}-\epsilon m_{1}m_{0}\big{]}.$
If this condition is not satisfied, it is optimal to set the standard deviation to its level $\bar{\sigma}$, i.e., to provide no information. In the other case, the interior solution is given by
$\sigma^{\ast}_{\infty}:=\frac{1}{2}\Bigg{[}\sqrt{\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Big{)}^{2}+4\frac{\delta-\epsilon m_{0}m_{1}\underline{\sigma}}{1+\epsilon^{2}m_{1}^{2}}}+\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Bigg{]},\quad\text{with}\quad\delta:=\frac{\eta}{c\Delta\sigma}.$
@@
(d) (e) (f)
Refer to caption Refer to caption Refer to caption
Figure 2: As a function of $\epsilon$ (a) $m^{\ast}_{\infty}$ (b) $\sigma^{\ast}_{\infty}$ (c) Damage cost with optimal information provision and without any information provision. Parameters: $\ell=a=1$, $\gamma=0.1$, $c=0.5$, $\lambda_{a}=0.75$, $\lambda_{q}=0.25$; $\kappa=0.05$ for (a), (b) and (c); $\kappa=0.09$ for (d), (e) and (f).
Figure 2 (a), (b), (c) and (d) display $m_{\infty}^{\ast}$ and $\sigma_{\infty}^{\ast}$ as a function of the best-in-class parameter $\epsilon$ in a situation where the firm puts more weight on its own personal target rather than the target of going close to the best-in-class target $q^{\ast}$. At equilibrium, only the product $\epsilon\lambda_{q}$ matters to measure the intensity with which the representative firm takes into account the best-in-class target. The main observation to be made on the information provision process is that it is a substitute to best-in-class objective. Indeed, we observe first that in (b), $\sigma_{\infty}^{\ast}$ increases with $\epsilon$, meaning that the higher the best-in-class objective intensity, the less the regulator is prone to impose stringent accounting rules. This phenomenon appears more clearly when the mean-reversion coefficient $\kappa$ is increased. In (a) and (b) its value is $\kappa=0.05$ while it is increased to $\kappa=0.09$ in (d) and (e). Increasing $\kappa$ leads to increasing the speed at which firms can identify the true value of their carbon footprint. We observe then that for large values of $\epsilon$, the information provision is no longer needed. This observation translates in terms of costs. Figure 2 (c) shows the decomposition of the cost function of the Sender. The black dotted line represents the cost in the absence of information provision, and thus, the damage. Providing information (figure (c) and (f)), always improves the objective function of the regulator whatever the best-in-class intensity $\epsilon$. However, the gain decreases with $\epsilon$ and might disappear in case of large mean-reversion values.
As a result, if we interpret the information provision as the level of stringency of the accounting rules, our model of information provision shows that the rules are a substitute for best-in-class objective target of a set of firms.
@@ Acknowledgements:
The authors thanks for their valuable comments in improving this manuscript the attendees of the Bachelier Congress 2024 edition, the Hammamet Conference on Stochastic Control and Games, the Banff BIRS workshop on “Modeling, Learning and Understanding”, the Online Seminar of International Bachelier Society, the Oxford seminar of applied mathematics, the Bachelier seminar at IHP, Paris, and in particular Agostino Capponi, Robert Crowell, Mike Ludkovski, Marcel Nutz, Fenghui Yu and Yufei Zhang.
@@ A.2 Carbon footprint accounting rules
The function in the Receiver’s objective functional is $f(x,v)=(\lambda_{q}+\lambda_{a})x^{2}-2(a\lambda_{a}+q\lambda_{q})x+\lambda_{q}q^{2}+\lambda_{a}a^{2}+\frac{1}{2\gamma}{v^{2}}$, with $\lambda_{a},\lambda_{q},\gamma>0$. Hence, we have $F_{0}=\lambda_{q}q^{2}+\lambda_{a}a^{2},F_{1}=-2a\lambda_{a}-2q\lambda_{q},F_{2}=\lambda_{q}+\lambda_{a},C_{1}=0,C_{2}=\frac{1}{2\gamma}$. The function in the Sender functional is $g(x,v)=\frac{1}{2}cx^{2}$, with $c>0$.
@@
Moreover, we provide hereafter a detailed version of the computations that leads to the solution to illustrate the easyness of the method.
We have $dX_{t}=\kappa(\ell-X_{t})dt+v_{t}dt+dW_{t}$ and $dM_{t}=bX_{t}dt+dB_{t}$ and the Receiver’s rate of profit is $f(x,v):=c(v)+\lambda(x-q)^{2}+\lambda_{a}(x-a)^{2}$ with $c(v):=\frac{1}{2}v^{2}/\gamma$. The estimated process dynamics is $dY_{t}=\kappa(\ell-Y_{t})dt+v_{t}dt+bP_{b}(t)dI_{t}$ where $I$ is the innovation process. The variance of the estimator $P_{b}(t)$ satisfies $\dot{P}_{b}=-2\kappa P_{b}-b^{2}P_{b}^{2}+1$, with $P_{b}(0)=0$, and we are only interested in its stationary value given by $-2\kappa P_{\infty}-b^{2}P_{\infty}^{2}+1=0$. The value function $V$ of the Receiver satisfies the HJB equation
$\zeta=\frac{1}{2}P_{\infty}^{2}V_{yy}+\inf_{v}\big{\{}k(\ell-y)V_{y}+vV_{y}+f(y,v)\},\quad P_{\infty}=\frac{1}{b^{2}}\sqrt{\kappa^{2}+b^{2}}-\frac{\kappa}{b^{2}}.$
From which we get $v^{\ast}=-\gamma V_{y}$ and thus
$\zeta=\frac{1}{2}P_{\infty}^{2}V_{yy}+\kappa(\ell-y)V_{y}-\frac{1}{2}\gamma V_{y}^{2}+\lambda_{a}(y-a)^{2}+\lambda_{q}(y-q)^{2}.$
Assuming $V(y)=Ay^{2}+By+C$ we get
$-2\kappa A-2\gamma A^{2}+\bar{\lambda}=0,\quad 2\kappa\ell A-\kappa B-2\gamma AB-2\bar{q}=0,\quad\bar{\lambda}:=\lambda_{a}+\lambda_{q},\quad\bar{q}:=a\lambda_{a}+q\lambda_{q}$
Hence, $V$ being convex we pick $A=\frac{\beta-\kappa}{2\gamma}$ $>0$, with $\beta:=\sqrt{\kappa^{2}+2\gamma\bar{\lambda}}$. And we get
$B=\frac{1}{\gamma\beta}\Big{(}\kappa\ell(\beta-\kappa)-2\gamma\bar{q}\Big{)}.$
Hence, the optimal control is
$v^{\ast}(y)=-2\gamma Ay-\gamma B=(\kappa-\beta)y-\frac{1}{\beta}\Big{(}\kappa\ell(\beta-\kappa)-2\gamma\bar{q}\Big{)}$
Thus the dynamics of the estimated process is
$d\hat{X}_{t}$ $=\kappa(\ell-\hat{X}_{t})dt+(\kappa-\beta)\hat{X}_{t}dt-\frac{1}{\beta}\Big{(}\kappa\ell(\beta-\kappa)-2\gamma\bar{q}\Big{)}dt+bP_{b}(t)dI_{t},$
$=-\beta\hat{X}_{t}dt+\frac{\ell\kappa^{2}+2\gamma\bar{q}}{\beta}dt+bP_{b}(t)dI_{t}=-\beta\Big{(}\hat{X}_{t}-\frac{\ell\kappa^{2}+2\gamma\bar{q}}{\beta^{2}}\Big{)}dt+bP_{b}(t)dI_{t}$
Hence, the dynamics of the estimated process is
$d\hat{X}_{t}$ $=\beta\big{(}\hat{\ell}(q)-\hat{X}_{t}\big{)}dt+bP_{b}(t)dI_{t},\quad\hat{\ell}(q):=\frac{\ell\kappa^{2}+2\gamma\bar{q}}{\kappa^{2}+2\gamma\bar{\lambda}}$
Using the fact that $dI_{t}=dM_{t}-b\hat{X}_{t}dt$, we have also
$d\hat{X}_{t}$ $=\big{[}\beta\big{(}\hat{\ell}(q)-\hat{X}_{t}\big{)}+b^{2}P_{b}(t)(X_{t}-\hat{X}_{t})\big{]}dt+bP_{b}(t)dB_{t},$
Besides, the dynamics of the state process $X$ is given by
$dX_{t}$ $=\big{[}\kappa(\ell-X_{t})+(\kappa-\beta)\hat{X}_{t}-\kappa\ell+\beta\hat{\ell}(q)\big{]}dt+dW_{t}$
$=\big{[}-\kappa X_{t}+(\kappa-\beta)\hat{X}_{t}+\beta\hat{\ell}(q)\big{]}dt+dW_{t}$
$=\big{[}\kappa(\hat{X}_{t}-X_{t})+\beta(\hat{\ell}(q)-\hat{X}_{t})\big{]}dt+dW_{t}$
We define $m_{t}:=\mathbb{E}(X_{t}^{*})$ and $U_{t}:=\mathbb{E}[{(X_{t}^{*})}^{2}]$, $Y_{t}:=\mathbb{E}[{(\widehat{X}_{t}^{*})}^{2}]$ and $Z_{t}:=\mathbb{E}[\widehat{X}_{t}^{*}X_{t}^{*}]$ and we will need the following:
$dm_{t}$ $=\beta(\hat{\ell}(q)-m_{t})dt,\quad dU_{t}=\big{[}2(\kappa-\beta)Z_{t}-2\kappa U_{t}+2\beta m_{t}\hat{\ell}(q)+1\big{]}dt,$
$dY_{t}$ $=\big{[}-2Y_{t}(\beta+b^{2}P_{b}(t))+2b^{2}P_{b}(t)Z_{t}+2\beta\hat{\ell}(q)m_{t}+b^{2}P_{b}^{2}(t)\big{]}dt,$
$dZ_{t}$ $=\big{[}\kappa(Y_{t}-Z_{t})+\beta(\hat{\ell}(q)m_{t}-Y_{t})+\beta(\hat{\ell}(q)m_{t}-Z_{t})+b^{2}P_{b}(t)(U_{t}-Z_{t})\big{]}dt.$
$dZ_{t}$ $=\big{[}(\kappa-\beta)Y_{t}-(\kappa+\beta+b^{2}P_{b}(t))Z_{t}+b^{2}P_{b}(t)U_{t}+2\beta\hat{\ell}(q)m_{t}\big{]}dt.$
Clearly $m_{\infty}=\hat{\ell}(q)$ and the stationary solutions of the above ODEs $U_{\infty},Y_{\infty},Z_{\infty}$ solve the system:
$\big{[}2(\kappa-\beta)Z_{\infty}-2\kappa U_{\infty}+2\beta(\hat{\ell}(q))^{2}+1\big{]}=0,$
$\big{[}-2Y_{\infty}(\beta+b^{2}P_{b}(\infty))+2b^{2}P_{b}(\infty)Z_{\infty}+2\beta(\hat{\ell}(q))^{2}+b^{2}P_{b}^{2}(\infty)\big{]}=0,$
$\big{[}\kappa(Y_{\infty}-Z_{\infty})+\beta((\hat{\ell}(q))^{2}-Y_{\infty})+\beta((\hat{\ell}(q))^{2}-Z_{\infty})+b^{2}P_{b}(\infty)(U_{\infty}-Z_{\infty})\big{]}=0$
The stationary variance is:
$\sigma^{2}_{\infty}(b)=U_{\infty}-(m_{\infty})^{2},$
hence, $X^{\ast}_{\infty}$ is Gaussian with mean ^$\hat{\ell}(q)$ and variance given by $\sigma^{2}_{\infty}(b)$.
Now, either we solve the above system or, we exploit the general formulas in Eq. (3.35) and (3.36), namely
$\mathrm{d}m_{t}$ $=\left(-\beta m_{t}+\frac{\kappa^{2}\ell+2\gamma\bar{q}}{\beta}\right)\mathrm{d}t,\quad m_{0}=x_{0}^{\top}$ (A.1)
$\mathrm{d}w_{t}$ $=(w_{t}\Theta(t)^{\top}+\Theta(t)w_{t}+\Xi(t)\Xi(t)^{\top})\mathrm{d}t,\quad w_{0}=\mathbf{0},$ (A.2)
with
$\Theta(t)$ $:=\begin{pmatrix}-\kappa&\kappa-\beta\\ P_{b}(t)b^{2}&-\beta-P_{b}(t)b^{2}\end{pmatrix}$
and
$\Xi(t):=\begin{pmatrix}1&0\\ 0&P_{b}(t)b\end{pmatrix}.$
So, we find $m_{\infty}=\frac{\ell\kappa^{2}+2\gamma\bar{q}}{\beta^{2}}=\hat{\ell}(q)$ as above and, introducing the moments
$w_{1}:=\mathbb{V}\text{\rm ar}(X_{\infty}^{*}),\qquad w_{2}=\mathbb{C}\text{\rm ov}(X_{\infty}^{*},{\widehat{X}_{\infty}}^{*}),\qquad w_{3}=\mathbb{V}\text{\rm ar}({\widehat{X}_{\infty}}^{*})$
we have that they solve the system
$-2\kappa w_{1}+2(\kappa-\beta)w_{2}+1=0,$
$P_{b}(\infty)b^{2}w_{1}-w_{2}(\beta+\kappa+P_{b}(\infty)b^{2})+(\kappa-\beta)w_{3}=0,$
$0\displaystyle 2P_{b}(\infty)b^{2}w_{2}+P_{b}^{2}(\infty)b^{2}-2w_{3}(\beta+P_{b}(\infty)b^{2})=$ 0 .
Recalling that $P_{b}(\infty)=\frac{1}{b^{2}}\big{(}\sqrt{\kappa^{2}+b^{2}}-\kappa\big{)}$, we find $w_{2}$
$w_{2}=\frac{1}{2\beta}\left(\sqrt{1+\frac{\kappa^{2}}{b^{2}}}-\frac{\kappa}{b}\right)^{2},$
and then our final target $\sigma_{b}^{2}(\infty)$:
$\quad w_{1}=\sigma_{b}^{2}(\infty)=\frac{1}{2\kappa}\left[1-\frac{\beta-\kappa}{\beta}\left(\sqrt{1+\frac{\kappa^{2}}{b^{2}}}-\frac{\kappa}{b}\right)^{2}\right]$
The fixed point condition for the equilibrium $q$ is given by $q=\mathbb{E}\big{[}X_{\infty}\big{]}-\epsilon\sigma_{\infty}(b)$ with $\mathbb{E}\big{[}X_{\infty}\big{]}=\hat{\ell}(q)$. Hence we have
$q=\frac{\ell\kappa^{2}+2\gamma\bar{q}}{\kappa^{2}+2\gamma\bar{\lambda}}-\epsilon\sigma_{\infty}(b),\quad\text{thus}\quad q^{\ast}(b):=\frac{\ell\kappa^{2}+2\gamma a\lambda_{a}}{\kappa^{2}+2\gamma\lambda_{a}}-\epsilon\frac{\kappa^{2}+2\gamma\bar{\lambda}}{\kappa^{2}+2\gamma\lambda_{a}}\sigma_{\infty}(b).$
Besides, the stationary equilibrium mean $m_{\infty}^{\ast}(b)$ is given by $m_{\infty}^{\ast}(b)=\hat{\ell}(q^{\ast}(b))$. Using the fact that ^$\hat{\ell}(q^{\ast}(b))=q^{\ast}(b)+\epsilon\sigma_{\infty}(b)$, we get
$m_{\infty}^{\ast}(b)$ $=\frac{\ell\kappa^{2}+2\gamma a\lambda_{a}}{\kappa^{2}+2\gamma\lambda_{a}}+\epsilon\sigma_{\infty}(b)\Big{[}1-\frac{\kappa^{2}+2\gamma\bar{\lambda}}{\kappa^{2}+2\gamma\lambda_{a}}\Big{]},$
$m_{\infty}^{\ast}(b)$ $=\frac{\ell\kappa^{2}+2\gamma a\lambda_{a}}{\kappa^{2}+2\gamma\lambda_{a}}-\epsilon\sigma_{\infty}(b)\frac{2\gamma\lambda_{q}}{\kappa^{2}+2\gamma\lambda_{a}}=:m_{0}-\epsilon m_{1}\sigma_{\infty}(b).$
Now, using $\mathbb{E}[(X^{\ast}_{\infty})^{2}]=(m_{\infty}^{\ast})^{2}+\sigma_{\infty}^{2}$, the optimisation problem of the Sender can be reformulated as
$\inf_{z\in(\underline{\sigma},\bar{\sigma}]}G(z)+\frac{1}{2}c\big{[}(m_{0}-\epsilon m_{1}z)^{2}+z^{2}\big{]},\quad G(z)=-\frac{\eta}{\Delta\sigma}\ln\frac{z-\underline{\sigma}}{\Delta\sigma},\quad z\in(\underline{\sigma},\bar{\sigma}],\quad\Delta\sigma=\overline{\sigma}-\underline{\sigma},$
Interior solution condition reads
$0\leq G^{\prime}(\overline{\sigma})+c\big{[}-\epsilon m_{1}(m_{0}-\epsilon m_{1}\overline{\sigma})+\overline{\sigma}\big{]},$
$\delta\leq\Delta\sigma\big{[}(1+\epsilon^{2}m_{1}^{2})\overline{\sigma}-\epsilon m_{1}m_{0}\big{]}$
with $\delta:=\eta/(c\Delta\sigma)$. Assuming the condition above, first-order optimality condition is
$-\frac{\eta}{\Delta\sigma(z-\underline{\sigma})}+c\big{[}(1+\epsilon^{2}m_{1}^{2})z-\epsilon m_{1}m_{0}\big{]}=0$
$-\eta+c\Delta\sigma(z-\underline{\sigma})\big{[}(1+\epsilon^{2}m_{1}^{2})z-\epsilon m_{1}m_{0}\big{]}=0$
$-\eta+c\Delta\sigma\big{[}(1+\epsilon^{2}m_{1}^{2})z^{2}-\big{(}\epsilon m_{0}m_{1}+(1+\epsilon^{2}m_{1}^{2})\underline{\sigma}\big{)}z+\epsilon m_{1}m_{0}\underline{\sigma}\big{]}=0$
We have
$(1+\epsilon^{2}m_{1}^{2})z^{2}-\big{(}\epsilon m_{0}m_{1}+(1+\epsilon^{2}m_{1}^{2})\underline{\sigma}\big{)}z+\epsilon m_{1}m_{0}\underline{\sigma}-\delta=0$
$z^{2}-\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Big{)}z-\frac{\delta-\epsilon m_{1}m_{0}\underline{\sigma}}{1+\epsilon^{2}m_{1}^{2}}=0$
$z^{\ast}$ $=\frac{1}{2}\Bigg{[}\sqrt{\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Big{)}^{2}+4\frac{\delta-\epsilon m_{1}m_{0}\underline{\sigma}}{1+\epsilon^{2}m_{1}^{2}}}+\Big{(}\frac{\epsilon m_{0}m_{1}}{1+\epsilon^{2}m_{1}^{2}}+\underline{\sigma}\Big{)}\Bigg{]}.$
$\Box$